# Self-Training Meets Consistency: Improving LLMs’ Reasoning With Consistency-Driven Rationale Evaluation

### v3

@@ Abstract
Jaehyeok Lee1, Keisuke Sakaguchi2, JinYeong Bak1 1Sungkyunkwan University 2Tohoku University hjl8708@skku.edu, keisuke.sakaguchi@tohoku.ac.jp, jy.bak@skku.edu
@@ 1 Introduction
In self-training approaches, accurately evaluating the quality of generated rationales is essential. Previous studies have evaluated rationale quality by examining whether the generated rationales lead to the correct answer to a given question Zelikman et al. (2022); Hoffman et al. (2023); Feng et al. (2024); Hosseini et al. (2024); Singh et al. (2024). However, using the correctness of a single prediction is unstable, as LLMs can reach correct answers through inappropriate reasoning steps Bao et al. (2024). Figure 1 shows an example of two generated rationales, $r^{1}$ and $r^{2}$. Despite $r^{2}$ shows incomplete reasoning, previous approaches would consider both rationales equally appropriate since they both lead to the correct answer for Q. Training models on such inappropriate rationales can cause them to learn flawed reasoning patterns.
To address this problem, we propose CREST (Consistency-driven Rationale Evaluation for Self-Training), a novel framework for LLM self-training. The core idea of CREST is to further evaluate rationales using follow-up questions that ask whether each answer option in the original question is correct or not. We first generate diverse rationales and evaluate them with an LLM as shown in Figure 1. Subsequently, we train the LLM on these rationales, rewarding rationales that lead to more consistent predictions (i.e., $r^{1}$) and penalizing those that lead to less consistent predictions (i.e., $r^{2}$). To achieve this, we propose two methods: rationale filtering and preference learning. In rationale filtering, we remove rationales that lead to incorrect answers in more than a certain number of follow-up questions during the supervised fine-tuning process. In preference learning, we train the model on mixed preferences from results of both original and follow-up questions, to favor rationales that result in correct answers in a greater number of follow-up questions.
@@ 2.1 Self-Training Approaches
The Chain-of-Thought (CoT) approach demonstrates that generating a step-by-step reasoning path before the final prediction enhances an LLM’s reasoning abilities Wei et al. (2022). Training LLMs on rationale data generated by humans Chung et al. (2024) or advanced models like GPT-4 further enhances reasoning abilities Liu et al. (2023). However, since high-quality rationale data is expensive to obtain, a number of approaches focus on training language models using self-generated rationales. STaR Zelikman et al. (2022), an early type of self-training approach, trains the language model by selecting the correct rationales based on binary feedback regarding the correctness of the answers generated by these rationales. RFT Yuan et al. (2023) enhances supervised data by generating and collecting diverse correct reasoning paths, focusing on mathematical reasoning. Other approaches, such as V-STaR, Iterative RPO, and Self-motivated Learning, also utilize incorrect rationales Feng et al. (2024); Hosseini et al. (2024); Pang et al. (2024) and adopt preference learning techniques, such as Proximal Policy Optimization (PPO) Schulman et al. (2017) and Direct Preference Optimization (DPO) Rafailov et al. (2023). Self-Explore Hwang et al. (2024) provides fine-grained rewards by identifying incorrect steps within the rationales. CREST provides fine-grained rewards through evaluating a rationale multiple times using follow-up questions augmented from the original dataset.
@@ 3.4 Rationale Evaluation
To generate follow-up questions which are closely related to the problem-solving process of each question in $\mathcal{D}$, we utilize the characteristics of multiple-choice questions: the solving process involves not only identifying the correct answer but also eliminating the incorrect options. We design each follow-up question to ask whether each of the answer options in the original question is correct or not. This type of follow-up question is used to evaluate robustness of reasoning ability in multiple-choice question-answering datasets Wang et al. (2024). Figure 1 shows an example of the follow-up questions and the evaluation.
@@ 3.6.1 Preference Pair Dataset Construction
We construct two preference pairs $P_{z}$ and $P_{\tilde{z}}$, which represent rationale preferences based on the rewards $z$ and $\tilde{z}$, respectively. We generate preference pairs where rationales with higher rewards $r^{w}$ are preferred over those with lower rewards $r^{l}$. Therefore, every pair consists of a question $q$, two generated rationales, and two predictions led by them: $(r^{w},p^{w})$ and $(r^{l},p^{l}$). We produce the final preference pair dataset $P_{\textit{total}}$ by combining $P_{z}$ and $P_{\tilde{z}}$ with $\lambda$ which satisfies $\lambda\in[0,1]$, ensuring that the proportion of $P_{\tilde{z}}$ in $P_{\textit{total}}$ is $\lambda$.
@@ 3.6.2 Training
We train $\textbf{{M}}_{\textbf{SFT}}$ on the preference pairs $P_{\textit{total}}$ using DPO, resulting in $\textbf{{M}}_{\textbf{CREST}}$. Given the preference pairs $P_{\textit{total}}$, the objective of this stage is to increase the log likelihood of preferred outputs over dispreferred ones:
@@ 4 Experiments
Table 1: Accuracy of various models across three reasoning datasets with Llama 3 8B and Gemma 7B model. ARC-C denotes challenge set in ARC test set. CREST consistently improves accuracy across all three datasets.
@@ 4.1 Experimental Settings
We conduct our experiments using the Llama 3 8B model AI@Meta (2024) and the Gemma 7B model Team et al. (2024) from HuggingFace Wolf et al. (2020), training them with Low-Rank Adaptation (LoRA) Hu et al. (2022).
@@
We generate rationales with temperature sampling with the following parameters: $T$=0.8, $TopP$=0.95, $N$=16, and max_new_tokens=512, then use greedy decoding for answer prediction. For supervised fine-tuning, we use epoch=6, batch size=32 and conduct learning rate search between {$65e-$ 6, $35e-$ 3}. For preference learning, we use $\beta$=0.1, epoch=4, batch size=8, and search max number of steps among {3000, 5000} and conduct learning rate search between {$75e-$ 7, $55e-$ 5} for all models. For more details about the prompts we use, please refer to Appendix E.
@@ 4.2 Baselines
Fine-tune (Label)
involves directly fine-tuning the base model on ground truth labels using a negative log-likelihood loss term, without relying on any generated rationales.
STaR Zelikman et al. (2022)
is an early approach for generating, filtering, and learning rationales using a generative language model. It generates a rationale for each question and trains the language model on rationales that lead to correct predictions. Additionally, STaR introduces a rationalization process that provides hints when the initial rationale fails to produce a correct prediction.
RFT Yuan et al. (2023)
stands for Rejection Sampling Fine-Tuning. RFT generates diverse rationales with a non-zero temperature and selects rationales to train based on binary feedback on the correctness of the final prediction. Unlike STaR, RFT does not have a rationalization process. In our experiments, $\textbf{{M}}_{\textbf{SFT}}$ with maximum tolerance corresponds to RFT.
Self-motivated Learning Feng et al. (2024)
exploits the inherent preference between correct rationales and incorrect rationales. It first trains a base model on generated and filtered rationales through supervised fine-tuning. It trains a reward model that assigns higher rewards to correct rationales than to incorrect ones. This reward model is then used to improve the reasoning performance of a supervised fine-tuned model through reinforcement learning using Proximal Policy Optimization (PPO) Schulman et al. (2017).
@@ 4.3 CREST
$\textit{M}_{\text{SFT}}$
is supervised fine-tuned on filtered rationales from the base model. The performance difference between this model and RFT demonstrates the effect of the rationale filtering process.
@@ 5.4 Evaluating Quality of Rationales
To qualitatively evaluate how the CREST impacts the model’s rationale generation, we randomly sample 100 questions from the ReClor validation set and evaluate the rationales from each model with GPT-4o. Following the methodology of Hwang et al. (2024), we use FLASK Ye et al. (2024) with GPT-4o to evaluate the logical robustness, correctness, and efficiency of the rationales. As shown in Table 2, CREST enhances rationale generation across all three metrics. Especially, rationale filtering in supervised fine-tuning improves the logical robustness and efficiency of the rationales. While preference learning on $P_{z}$ makes $\textbf{{M}}_{\textbf{SFT}}$ generate more logically correct rationales, it decreases the robustness of the rationales. However, preference learning on $P_{\textit{total}}$ yields higher performance across all three metrics compared to using only $P_{z}$. These evaluation results show that $\textbf{{M}}_{\textbf{CREST}}$ generates more logically robust and correct rationales than the baselines.
@@ 5.5 Evaluating CREST Models on Follow-up Questions
We evaluate the rationales generated by each trained model for the original questions in the ReClor validation set using follow-up questions, which is shown in Figure 6. As in the Rationale Generation and Evaluation stage, we input the generated rationales and follow-up questions into the base model (Llama 3 8B), then measure accuracy over all follow-up questions. To assess how different training methods affect the rationale generation, we adopt Zero-shot-CoT Kojima et al. (2022) as a baseline model. The improvement between RFT and $\textbf{{M}}_{\textbf{SFT}}$ shows the effect of rationale filtering in generating rationales that are more robust to follow-up questions. As shown in Figure 6, CREST trains the LLM to generate rationales that are more robust to follow-up questions.
@@ 8 Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2024-00509258, Global AI Frontier Lab) and abductive inference framework using omni-data for understanding complex causal relations & ICT Creative Consilience program (2022-0-00680).
Appendix A Phi-2 Experiment Result
@@
To demonstrate the robustness of CREST, we also test CREST with Phi-2 model Javaheripi et al. (2023). Phi-2 has 2.7B parameters, which is much smaller compared to Llama 3 8B and Gemma 7B which have 8.0B and 8.5B parameters, respectively. As shown in Table 3, CREST outperforms other self-training baselines across the three datasets, and preference learning to Fine-tune (Label) model consistently improves performance. This result shows that CREST can function effectively with this relatively small model.
A. freeway B. radio C. wires D. computer network E. electrical circuit
$\tilde{z}=5$
$z=1$
$\tilde{z}=4$
$z=1$
$\tilde{z}=3$
$z=1$
$\tilde{z}=2$
$z=1$
Appendix B Data and Rationale Statistics
@@
Appendix C Rationale Generation and Evaluation Case Study
@@
Appendix D Detailed Preference Pair Datasets Construction
This section presents a more detailed algorithm for constructing the preference pair dataset used in preference learning. As shown in Algorithm 1, we construct two preference pair sets, $P_{z}$ and $P_{\tilde{z}}$, based on $z$ and $\tilde{z}$, respectively. Then, we construct $P_{\text{total}}$ such that the proportion of $P_{\tilde{z}}$ in $P_{\text{total}}$ is $\lambda$, and the proportion of $P_{z}$ is $1-\lambda$, by randomly sampling pairs from $P_{z}$ and $P_{\tilde{z}}$ and combining them.
@@ Appendix D Detailed Preference Pair Datasets Construction
Appendix E Prompts
In this section, we introduce the prompt templates used for rationale generation and inference. We construct input text for the language model based on these templates. All the prompt templates we present are designed for the ReClor dataset Yu et al. (2020). Unlike ReClor, the ARC Clark et al. (2018) and CSQA Talmor et al. (2019) datasets do not include a passage, so we use different prompt templates for them. As a result, the [Question] part in the prompt templates for ARC and CSQA consists only of the question and the answer choices.
[Instruction]

### v4

@@ Abstract
[…]
**footnotetext: Corresponding authors
@@ 1 Introduction
In self-training approaches, accurately evaluating the quality of generated rationales is essential. Previous studies have evaluated rationale quality by examining whether the generated rationales lead to the correct answer to a given question Zelikman et al. (2022); Hoffman et al. (2023); Feng et al. (2024); Hosseini et al. (2024); Singh et al. (2024). However, using the correctness of a single prediction is unstable, as LLMs can reach correct answers through inappropriate reasoning steps Bao et al. (2024a). Figure 1 shows an example of two generated rationales, $r^{1}$ and $r^{2}$. Despite $r^{2}$ shows incomplete reasoning, previous approaches would consider both rationales equally appropriate since they both lead to the correct answer for Q. Training models on such inappropriate rationales can cause them to learn flawed reasoning patterns.
To address this problem, we propose CREST (Consistency-driven Rationale Evaluation for Self-Training), a novel framework for LLM self-training. The core idea of CREST is to further evaluate rationales using follow-up questions that ask whether each answer option in the original question is correct or not. We first generate diverse rationales using temperature sampling and evaluate them with an LLM as shown in Figure 1. Subsequently, we train the LLM on these rationales, rewarding rationales that lead to more consistent predictions (i.e., $r^{1}$) and penalizing those that lead to less consistent predictions (i.e., $r^{2}$). To achieve this, we propose two methods: rationale filtering and preference learning. In rationale filtering, we remove rationales that lead to incorrect answers in more than a certain number of follow-up questions during the supervised fine-tuning process. In preference learning, we train the model on mixed preferences from results of both original and follow-up questions, to favor rationales that result in correct answers in a greater number of follow-up questions.
@@ 2.1 Self-Training Approaches
Chain-of-Thought (CoT) approach demonstrates that generating a step-by-step reasoning path before the final prediction enhances an LLM’s reasoning abilities Wei et al. (2022). Training LLMs on rationale data generated by humans Chung et al. (2024) or advanced models like GPT-4 further enhances reasoning abilities Liu et al. (2023). However, since high-quality rationale data is expensive to obtain, a number of approaches focus on training language models using self-generated rationales. STaR Zelikman et al. (2022), an early type of self-training approach, trains the language model by selecting the correct rationales based on binary feedback regarding the correctness of the answers generated by these rationales. RFT Yuan et al. (2023) enhances supervised data by generating and collecting diverse correct reasoning paths, focusing on mathematical reasoning. Other approaches, such as V-STaR, Iterative RPO, and Self-motivated Learning, also utilize incorrect rationales Feng et al. (2024); Hosseini et al. (2024); Pang et al. (2024) and adopt preference learning techniques, such as Proximal Policy Optimization (PPO) Schulman et al. (2017) and Direct Preference Optimization (DPO) Rafailov et al. (2023). Self-Explore Hwang et al. (2024) provides fine-grained rewards by identifying incorrect steps within the rationales. Wei Jie et al. (2024) proposes a self-training framework that exposes a model to each question multiple times using temperature sampling, thereby assessing the model’s confidence in the given question. CREST provides fine-grained rewards through evaluating a self-training rationale multiple times using follow-up questions augmented from the original dataset, emphasizing the rationale’s ability to consistently lead to correct answers.
@@ 3.4 Rationale Evaluation
To generate follow-up questions that are closely related to the problem-solving process of each question in $\mathcal{D}$, we utilize the characteristics of multiple-choice questions: the solving process involves not only identifying the correct answer but also eliminating the incorrect options. We design each follow-up question to ask whether each of the answer options in the original question is correct or not. This type of follow-up question is used to evaluate the robustness of reasoning ability in multiple-choice question-answering datasets Wang et al. (2024). Figure 1 shows an example of the follow-up questions and the evaluation.
@@ 3.6.1 Preference Pair Dataset Construction
We construct the preference pair dataset $P_{\textit{total}}$ for preference learning by first creating two sets of preference pairs $P_{z}$ and $P_{\tilde{z}}$, which represent rationale preferences based on the rewards $z$ and $\tilde{z}$, respectively. $P_{\textit{total}}$ is then formed by randomly sampling pairs from these two sets.
To construct $P_{z}$ and $P_{\tilde{z}}$, we generate preference pairs in which rationales with higher rewards $r^{w}$ are preferred over those with lower rewards $r^{l}$, based on $z$ and $\tilde{z}$, respectively. Each preference pair consists of a question $q$, two generated rationales, and their corresponding predictions $p^{w}$ and $p^{l}$: $(q,r^{w},p^{w},r^{l},p^{l}$). Algorithm 1 outlines the detailed procedure for generating $P_{z}$ and $P_{\tilde{z}}$.
Then, we construct $P_{\textit{total}}$ by sampling pairs from $P_{z}$ and $P_{\tilde{z}}$ with a weighting factor $\lambda$, which controls the relative contribution of rationale preferences derived from $z$ and $\tilde{z}$ during preference learning. The parameter $\lambda$ satisfies $\lambda\in[0,1]$, ensuring that the proportion of $P_{\tilde{z}}$ in $P_{\textit{total}}$ is $\lambda$. For instance, if a total of 10,000 pairs are used for preference learning and $\lambda=0.4$, $P_{\textit{total}}$ would consist of 4,000 randomly selected pairs from $P_{\tilde{z}}$ and 6,000 randomly selected pairs from $P_{z}$. The total number of pairs used for preference learning is determined by the maximum number of training steps multiplied by the batch size.
@@ 3.6.2 Training
We train $\textbf{{M}}_{\textbf{SFT}}$ on the preference pairs $P_{\textit{total}}$ using DPO, resulting in $\textbf{{M}}_{\textbf{CREST}}$. Given the preference pairs $P_{\textit{total}}$, the objective of this stage is to increase the log-likelihood of preferred outputs over dispreferred ones:
@@ 4 Experiments
Table 1: Accuracy of various models across three reasoning datasets with Llama 3 8B and Gemma 7B model. ARC-C denotes the challenge set in the ARC test set. CREST consistently improves accuracy across all three datasets.
@@ 4.1 Experimental Settings
We conduct our experiments using the Llama 3 8B model [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B AI@Meta (2024) and the Gemma 7B model [^3] https://huggingface.co/google/gemma-7b Team et al. (2024) from HuggingFace Wolf et al. (2020), training them with Low-Rank Adaptation (LoRA) Hu et al. (2022).
@@
We generate rationales with temperature sampling with the following parameters: $T$=0.8, $TopP$=0.95, $N$=16, and max_new_tokens=512, then use greedy decoding for answer prediction. For supervised fine-tuning, we use epoch=6, batch size=32 and conduct learning rate search between {$65e-$ 6, $35e-$ 3}. For preference learning, we use $\beta$=0.1, epoch=4, batch size=8, and search max number of steps among {3000, 5000} and conduct learning rate search between {$75e-$ 7, $55e-$ 5} for all models. The input and output prompt templates for model evaluation are illustrated in Figures 7 and 10. For more details about the prompts used in this study, please refer to Appendix F.
@@ 4.2 Baselines
  • Fine-tune (Label) involves directly fine-tuning the base model on ground truth labels using a negative log-likelihood loss term, without relying on any generated rationales.
  • STaR Zelikman et al. (2022) is an early approach for generating, filtering, and learning rationales using a generative language model. It generates a rationale for each question and trains the language model on rationales that lead to correct predictions. Additionally, STaR introduces a rationalization process that provides hints when the initial rationale fails to produce a correct prediction.
  • RFT Yuan et al. (2023) stands for Rejection Sampling Fine-Tuning. RFT generates diverse rationales with a non-zero temperature and selects rationales to train based on binary feedback on the correctness of the final prediction. Unlike STaR, RFT does not have a rationalization process. In our experiments, $\textbf{{M}}_{\textbf{SFT}}$ with maximum tolerance corresponds to RFT.
  • Self-motivated Learning Feng et al. (2024) exploits the inherent preference between correct rationales and incorrect rationales. It first trains a base model on generated and filtered rationales through supervised fine-tuning. It trains a reward model that assigns higher rewards to correct rationales than to incorrect ones. This reward model is then used to improve the reasoning performance of a supervised fine-tuned model through reinforcement learning using Proximal Policy Optimization (PPO) Schulman et al. (2017).
@@ 4.3 CREST
  • $\textit{M}_{\text{SFT}}$ is supervised fine-tuned on filtered rationales from the base model. The performance difference between this model and RFT demonstrates the effect of the rationale filtering process.
@@ 5.4 Evaluating Quality of Rationales
To qualitatively evaluate how the CREST impacts the model’s rationale generation, we randomly sample 100 questions from the ReClor validation set and evaluate the rationales from each model with GPT-4o. Following the methodology of Hwang et al. (2024), we employ FLASK Ye et al. (2024), a fine-grained evaluation protocol for model-based evaluation, which exhibits a high correlation with human-based evaluation. Specifically, we focus on the ‘logical thinking’ category in FLASK, which encompasses three aspects: logical correctness, logical robustness, and logical efficiency. Logical correctness evaluates the model’s ability to produce logically correct final answers. Logical robustness evaluates the generalizability of the step-by-step reasoning process without contradictions. Logical efficiency examines whether the reasoning process is concise and free of unnecessary steps. For the exact prompt templates used in the FLASK evaluation, please refer to Figures 11 and 12.
As shown in Table 2, CREST enhances rationale generation across all three aspects. Especially, rationale filtering in supervised fine-tuning improves the logical robustness and efficiency of the rationales. While preference learning on $P_{z}$ makes $\textbf{{M}}_{\textbf{SFT}}$ generate more logically correct rationales, it decreases the robustness of the rationales. However, preference learning on $P_{\textit{total}}$ yields higher performance across all three metrics compared to using only $P_{z}$. These evaluation results show that $\textbf{{M}}_{\textbf{CREST}}$ generates more logically robust and correct rationales than the baselines.
@@ 5.5 Evaluating CREST Models on Follow-up Questions
We evaluate the rationales generated by each trained model for the original questions in the ReClor validation set using follow-up questions, which is shown in Figure 6. As in the Rationale Generation and Evaluation stage, we input the generated rationales and follow-up questions into the base model (Llama 3 8B), then measure accuracy over all follow-up questions. To assess how different training methods affect the rationale generation, we employ Zero-shot-CoT Kojima et al. (2022) as a baseline model. The improvement between RFT and $\textbf{{M}}_{\textbf{SFT}}$ shows the effect of rationale filtering in generating rationales that are more robust to follow-up questions. As shown in Figure 6, CREST trains the LLM to generate rationales that are more robust to follow-up questions.
@@ 8 Acknowledgements
We would like to thank the anonymous reviewers for their helpful questions and comments. This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2024-00509258, Global AI Frontier Lab, No. RS-2024-00398115, Research on the reliability and coherence of outcomes produced by Generative AI and No. RS-2019-II190421, AI Graduate School Support Program(Sungkyunkwan University)) and abductive inference framework using omni-data for understanding complex causal relations & ICT Creative Consilience program (2022-0-00680).
Appendix A Phi-2 Experiment Results
@@
To demonstrate the robustness of CREST, we also test CREST with Phi-2 model [^4] https://huggingface.co/microsoft/phi-2 Javaheripi et al. (2023). Phi-2 has 2.7B parameters, which is much smaller compared to Llama 3 8B and Gemma 7B which have 8.0B and 8.5B parameters, respectively. As shown in Table 3, CREST outperforms other self-training baselines across the three datasets, and preference learning to Fine-tune (Label) model consistently improves performance. This result shows that CREST can function effectively with this relatively small model.
Appendix B Evaluating CoT Performance in Zero- and Few-Shot Settings
Base model Model ReClor ARC-C CSQA
Llama 3 8B Zero-shot 52.10 69.28 53.89
Zero-shot-CoT 45.70 63.74 48.57
Few-shot 55.30 77.21 70.76
Few-shot-CoT 40.60 74.23 71.25
Gemma 7B Zero-shot 53.60 77.47 65.68
Zero-shot-CoT 43.50 65.96 48.24
Few-shot 58.70 83.11 75.02
Few-shot-CoT 51.90 80.29 72.89
Phi-2 Zero-shot 42.00 68.51 57.74
Zero-shot-CoT 38.40 66.01 52.17
Few-shot 43.10 71.84 62.41
Few-shot-CoT 43.00 75.68 70.02
Table 4: Accuracy of base models in zero-shot and few-shot settings, with and without CoT prompting, on the three reasoning datasets. In many cases, CoT prompting results in performance degradation.
To measure the accuracy of M itself using Chain-of-Thoughts (CoT) without fine-tuning, we conduct experiments with M. Specifically, we examine the performance of $M$ instructed to generate the rationale and prediction, represented as $(r,p)=\textbf{{M}}(q)$. We refer to these approaches as Zero-shot-CoT (instructed to generate a rationale and prediction without examples) and Few-shot-CoT (given few-shot examples and then instructed to generate a rationale and prediction). The input-output format used for those CoT models is the same as the input-output format of $\textbf{{M}},\textbf{{M}}_{\textbf{SFT}},\textbf{{M}}_{\textbf{CREST}}$.
As shown in Table 4, these CoT approaches underperformed compared to their non-CoT counterparts in many cases. Some previous studies support this performance degradation. Wei et al. (2022) show that models with size not big enough would not benefit from chain-of-thought reasoning. Some studies Bao et al. (2024b); Xu et al. (2023) have reported common performance degradation with CoT approaches in complex reasoning tasks.
Appendix C Data and Rationale Statistics
@@
Appendix D Rationale Generation and Evaluation Case Study
A. freeway B. radio C. wires D. computer network E. electrical circuit
$\tilde{z}=5$
$z=1$
$\tilde{z}=4$
$z=1$
$\tilde{z}=3$
$z=1$
$\tilde{z}=2$
$z=1$
@@ Appendix D Rationale Generation and Evaluation Case Study
Appendix E Preference Pair Datasets Construction Algorithm
This section presents a more detailed algorithm for constructing the preference pair dataset used in preference learning. As shown in Algorithm 1, we construct two preference pair sets, $P_{z}$ and $P_{\tilde{z}}$, based on $z$ and $\tilde{z}$, respectively.
@@ Appendix E Preference Pair Datasets Construction Algorithm
Appendix F Prompts
In this section, we introduce the prompt templates used for rationale generation, inference, and evaluation with FLASK. We construct input text for the language model based on these templates. All the prompt templates we present are designed for the ReClor dataset Yu et al. (2020). Unlike ReClor, the ARC Clark et al. (2018) and CSQA Talmor et al. (2019) datasets do not include a passage, so we use different prompt templates for them. As a result, the [Question] part in the prompt templates for ARC and CSQA consists only of the question and the answer choices.
Input  [Instruction]